# Mixture of Experts in Large Language Models thanks:

### v1

@@ Abstract
Index Terms:
Large language models, mixture of experts, expert routing, meta learning, knowledge transfer, sparse activation, large language models architecture, natural language processing
@@ I Introduction and Fundamentals
Open-source initiatives have accelerated MoE adoption by releasing advanced models such as Jamba, Qwen1.5-MoE, and Mistral-8x22B. Ongoing improvements to routing and load balancing in Qwen3-MoE, Claude 3.5 Sonnet, and Llama 4 show that innovation now centers on efficiency, controllability, and multi-task generalization rather than scale alone.
@@
Param-Efficient LoRA-MoE (4) 2411.18466 / Nexus (8) 2408.15901 Frozen router 2023–24 $<$1 % parameter update (PEFT); “upcycles” dense checkpoints to adaptive MoE.
@@
for tree= child anchor=west, parent anchor=east, grow’=east, anchor=west, base=left, font=, rectangle, draw=black, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, language/.style=fill=softblue, multimodal/.style=fill=mintgreen, architecture/.style=fill=warmgray, training/.style=fill=lavender, routing/.style=fill=peach, application/.style=fill=sage, challenge/.style=fill=dustyrose, , where level=1text width=11em,font=,, where level=2text width=12.5em,font=,, where level=3text width=12em,font=,, where level=4text width=17em,font=,, [ Mixture of Experts (MoE), ver [  Language Models, language [  General LLMs, language [  Open Source: Mixtral, Qwen2-MoE, DeepSeek-MoE, Switch-T , language, text width=28em ] [  Commercial: GPT-4 (rumored), PaLM-2, GLaM, DBRX , language, text width=28em ] [  Research: GShard, ST-MoE, Expert Router , language, text width=28em ] ] [  Specialized LLMs, language [  Translation: NLLB-200, NLLB, Multilingual-MoE , language, text width=28em ] [  Code Generation: DeepSeek-Coder-MoE, CodeT5-MoE , language, text width=28em ] [  Domain-specific: MoGU, BioBERT-MoE , language, text width=28em ] ] ] [  Multimodal Models, multimodal [  Vision-Language, multimodal [  Understanding: LIMoE, MoE-LLaVA, LLaVA-MoLE , multimodal, text width=28em ] [  Generation: Omni-SMoLA, DALLE-MoE , multimodal, text width=28em ] ] [  Computer Vision, multimodal [  Image Classification: V-MoE, Swin-MoE, pMoE , multimodal, text width=28em ] [  Adversarial Robustness: ADVMOE , multimodal, text width=28em ] ] [  Multi-task Unified, multimodal [  Unified Models: Uni-MoE, MM1, MoCLE , multimodal, text width=28em ] ] ] [  Architectural Innovations, architecture [  Expert Selection, architecture [  Top-K Routing: Standard Top-K, Adaptive Top-K , architecture, text width=28em ] [  Soft Routing: Weighted Routing, Probabilistic Selection , architecture, text width=28em ] ] [  Structural Variants, architecture [  Hierarchical MoE: HMoE, Nested MoE , architecture, text width=28em ] [  Orthogonal MoE: OMoE , architecture, text width=28em ] [  Calibrated MoE: MoCaE , architecture, text width=28em ] ] [  Optimization Techniques, architecture [  Load Balancing: Auxiliary Loss, Expert Regularization , architecture, text width=28em ] [  Efficiency: DeepSpeed-MoE, FasterMoE, HyperMoE , architecture, text width=28em ] ] ] [  Training Strategies, training [  Expert Initialization, training [  Random Initialization, Pre-trained Expert Initialization , training, text width=28em ] ] [  Training Paradigms, training [  End-to-End Training, Stage-wise Training , training, text width=28em ] ] [  Regularization Methods, training [  Expert Dropout, Noise Injection, Sparsity Regularization , training, text width=28em ] ] ] [  Routing Mechanisms, routing [  Gating Networks, routing [  Learned Gating: Neural Gating, Attention-based Gating , routing, text width=28em ] [  Hash-based Routing: Fixed Hash, Learnable Hash , routing, text width=28em ] ] [  Load Balancing, routing [  Auxiliary Loss Functions, Expert Capacity Constraints , routing, text width=28em ] ] [  Dynamic Routing, routing [  Adaptive Expert Selection, Context-aware Routing , routing, text width=28em ] ] ] [  Application Scenarios, application [  Natural Language Processing, application [  Text Generation, Machine Translation, Question Answering , application, text width=28em ] ] [  Computer Vision, application [  Image Classification, Object Detection, Segmentation , application, text width=28em ] ] [  Multimodal Tasks, application [  Vision-Language Understanding, Cross-modal Retrieval , application, text width=28em ] ] [  Scientific Computing, application [  Drug Discovery, Climate Modeling, Protein Folding , application, text width=28em ] ] ] [  Challenges & Outlook, challenge [  Scalability Issues, challenge [  Expert Scaling, Memory Consumption, Communication Overhead , challenge, text width=28em ] ] [  Training Challenges, challenge [  Expert Collapse, Load Imbalance, Gradient Synchronization , challenge, text width=28em ] ] [  Deployment Challenges, challenge [  Inference Latency, Model Compression, Hardware Efficiency , challenge, text width=28em ] ] [  Future Directions, challenge [  Automated Expert Design, Continual Learning, Federated MoE , challenge, text width=28em ] ] ] ]
@@ II-A Foundational MoE Architectures
where TopK masks all but the $k$ largest scores with $\infty$- ∞ before applying softmax. This Noisy Top-$k$ routing mechanism shazeer2017outrageously promotes exploration, mitigates early expert collapse, and ensures sparse yet balanced expert activation.
@@
Figure 3: A brief illustration of sparsely gated Mixture of Experts (MoE) architecture on decoder only transformer. In this figure, the top-$k$ routing mechanism is configured with $k$=2, meaning the gating function selects the two highest-scoring FFN experts for each token based on the router’s softmax probabilities. The selected experts are evaluated in parallel, with their outputs aggregated using weighted summation.
@@
While MoEs offer high capacity through modular specialization, their lack of shared inductive bias can hinder generalization when routing is unstable. Training dynamics are sensitive to expert overlap and gate smoothness, and unlike Bayesian ensembles, MoEs delegate via hard routing rather than aggregate over uncertainty—potentially increasing variance. Nonetheless, large-scale evaluations show that MoEs achieve comparable performance to dense models on language tasks while activating $\sim$10$\times$ fewer parameters per token 2501.16352 , demonstrating an effective trade-off between compute efficiency and representational specialization.
@@ II-B Advanced Architectural Variants
where $\Delta\theta_{\text{train}}$ denotes updated parameters and $\theta_{\text{full}}$ is the total parameter set. Methods such as MoCE-IR 2411.18466 and Adamix wang2022adamix implement this by freezing shared layers and updating only expert heads or low-rank adapters, often achieving $<1\%$ parameter updates with minimal performance drop.
@@ IV-A Recommendation Systems and Search
In large-scale recommendation and search applications, MoE architectures are increasingly employed to handle the inherent complexity of multi-domain, multi-task personalization 2404.18465 ; hou2022towards . Conventional models often struggle to balance domain-specific signal extraction and shared knowledge transfer, particularly under varying user contexts and item scenarios.
To address these limitations, M3oE 2404.18465 proposes a modular MoE framework that jointly models domain- and task-level heterogeneity. It deploys three parallel expert modules to capture shared user preferences, domain-specific behavior, and task-specific patterns respectively. These modules are fused through a hierarchical gating design that allows adaptive aggregation of relevant signals across heterogeneous learning objectives. By integrating AutoML techniques for structure search, M3oE dynamically adjusts its expert composition, improving alignment with evolving personalization targets and ensuring performance scalability under real-world constraints.
As illustrated in Figure 6, AESM2 begins with a shared embedding lookup over input features, scenario indicators, and task indicators. The encoded representations are processed through stacked multi-scenario layers, where each layer refines contextual understanding across traffic segments or environments. On top of these layers, scenario-informed representations are passed into multi-task-specific towers to generate differentiated outputs. The architecture employs a hierarchical routing mechanism that automatically selects experts at both the scenario and task levels. It does so without manual intervention, allowing the model to allocate computation to scenario-shared or scenario-specialized experts as needed. This structure enables explicit control over knowledge transfer across scenarios and tasks while maintaining model flexibility. Evaluations show that AESM2 consistently improves retrieval quality and training stability under dynamic traffic distributions, outperforming static multi-gate MoE variants that lack automatic structure adaptation.

### v2

@@ I Introduction and Fundamentals
Open-source initiatives have accelerated MoE adoption through the release of advanced models such as Jamba, Qwen1.5-MoE, and Mistral-8x22B, while continued improvements to routing and load balancing in Qwen3-MoE, Claude 3.5 Sonnet, and Llama 4 indicate that innovation now centers on efficiency, controllability, and multi-task generalization rather than scale alone. From mid-2025 onward, the main technical focus further shifts from increasing parameter count to making routing reliable under long training runs and deployment constraints. On the routing side, Omi et al. omi2025similarity_router propose a similarity-preserving load-balancing objective that stabilizes expert selection for related inputs while avoiding expert collapse, enabling faster and less redundant training. Complementarily, Dong et al. dong2025maxscore formulate routing as a constrained optimization problem via MaxScore, addressing token dropping and padding inefficiency caused by hard expert-capacity constraints. At the survey level, Zhang et al. zhang2025moe_review summarize mid-2025 MoE design choices, including gating strategies, sparse and hierarchical variants, multimodal extensions, and deployment considerations, showing that evaluation increasingly emphasizes expert diversity, calibration, and inference aggregation. In late 2025, Qwen3-VL bai2025qwen3vl reports vision–language MoE variants (for example, 30B-A3B and 235B-A22B) designed for latency–quality trade-offs with long interleaved multimodal context, indicating that MoE is now treated as a practical control for compute allocation across modalities rather than a scaling-only mechanism.
@@
Param-Efficient LoRA-MoE (4) 2411.18466 / Nexus (8) 2408.15901 Frozen router 2023–24 <<1 % parameter update (PEFT); “upcycles” dense checkpoints to adaptive MoE.
@@
Recent work places stronger emphasis on deployment-aware evaluation and design. A comprehensive survey by Zhang et al. zhang2025moe_review highlights that modern assessment of MoE systems increasingly focuses on expert utilization balance, calibration, and inference-time aggregation behavior rather than parameter count. In parallel, systems such as Qwen3-VL bai2025qwen3vl demonstrate that routing policies can be shaped to meet practical latency and memory constraints in multimodal inference. Together, these results indicate a shift from MoE architectures optimized mainly for training scale toward designs shaped by robustness, efficiency, and deployment feasibility.
@@
Language Models General LLMs Open Source: Mixtral, Qwen2-MoE, DeepSeek-MoE, Switch-T Commercial: GPT-4 (rumored), PaLM-2, GLaM, DBRX Research: GShard, ST-MoE, Expert Router Specialized LLMs Translation: NLLB-200, NLLB, Multilingual-MoE Code Generation: DeepSeek-Coder-MoE, CodeT5-MoE Domain-specific: MoGU, BioBERT-MoE Multimodal Models Vision-Language Understanding: LIMoE, MoE-LLaVA, LLaVA-MoLE Generation: Omni-SMoLA, DALLE-MoE Computer Vision Image Classification: V-MoE, Swin-MoE, pMoE Adversarial Robustness: ADVMOE Multi-task Unified Unified Models: Uni-MoE, MM1, MoCLE Architectural Innovations Expert Selection Top-K Routing: Standard Top-K, Adaptive Top-K Soft Routing: Weighted Routing, Probabilistic Selection Structural Variants Hierarchical MoE: HMoE, Nested MoE Orthogonal MoE: OMoE Calibrated MoE: MoCaE Optimization Techniques Load Balancing: Auxiliary Loss, Expert Regularization Efficiency: DeepSpeed-MoE, FasterMoE, HyperMoE Training Strategies Expert Initialization Random Initialization, Pre-trained Expert Initialization Training Paradigms End-to-End Training, Stage-wise Training Regularization Methods Expert Dropout, Noise Injection, Sparsity Regularization Routing Mechanisms Gating Networks Learned Gating: Neural Gating, Attention-based Gating Hash-based Routing: Fixed Hash, Learnable Hash Load Balancing Auxiliary Loss Functions, Expert Capacity Constraints Dynamic Routing Adaptive Expert Selection, Context-aware Routing Application Scenarios Natural Language Processing Text Generation, Machine Translation, Question Answering Computer Vision Image Classification, Object Detection, Segmentation Multimodal Tasks Vision-Language Understanding, Cross-modal Retrieval Scientific Computing Drug Discovery, Climate Modeling, Protein Folding Challenges & Outlook Scalability Issues Expert Scaling, Memory Consumption, Communication Overhead Training Challenges Expert Collapse, Load Imbalance, Gradient Synchronization Deployment Challenges Inference Latency, Model Compression, Hardware Efficiency Future Directions Automated Expert Design, Continual Learning, Federated MoE
@@ II-A Foundational MoE Architectures
where TopK masks all but the $k$ largest scores with -$\infty$ before applying softmax. This Noisy Top-$k$ routing mechanism shazeer2017outrageously promotes exploration, mitigates early expert collapse, and ensures sparse yet balanced expert activation.
@@
Figure 3: A brief illustration of sparsely gated Mixture of Experts (MoE) architecture on decoder only transformer. In this figure, the top-$k$ routing mechanism is configured with kk=2, meaning the gating function selects the two highest-scoring FFN experts for each token based on the router’s softmax probabilities. The selected experts are evaluated in parallel, with their outputs aggregated using weighted summation.
@@
[…]
@@ II-B Advanced Architectural Variants
where $\Delta\theta_{\text{train}}$ denotes updated parameters and $\theta_{\text{full}}$ is the total parameter set. Methods such as MoCE-IR 2411.18466 and Adamix wang2022adamix implement this by freezing shared layers and updating only expert heads or low-rank adapters, often achieving $1\%$ parameter updates with minimal performance drop.
@@ IV-A Recommendation Systems and Search
In large-scale recommendation and search applications, MoE architectures are increasingly employed to handle the inherent complexity of multi-domain and multi-task personalization 2404.18465; hou2022towards. Conventional dense models often struggle to balance domain-specific signal extraction and shared knowledge transfer, especially under rapidly changing user contexts, traffic distributions, and item lifecycles.
To address these limitations, M3oE 2404.18465 proposes a modular MoE framework that jointly models domain- and task-level heterogeneity. It deploys parallel expert modules to capture shared user preferences, domain-specific behavior, and task-specific patterns, which are fused through a hierarchical gating mechanism. By integrating AutoML-based structure search, M3oE can adapt its expert composition over time, improving robustness and scalability under production workloads.
As illustrated in Figure 6, AESM2 follows a similar design philosophy but targets scenario-aware recommendation. It begins with a shared embedding layer over input features, scenario indicators, and task indicators, followed by stacked multi-scenario layers that refine contextual representations across traffic segments. Scenario-informed representations are then routed into task-specific towers through a hierarchical routing mechanism that selects experts at both scenario and task levels. This design enables controlled knowledge sharing while allowing specialization when traffic patterns diverge. Empirical results show that AESM2 improves retrieval quality and training stability under dynamic traffic shifts, outperforming static multi-gate MoE baselines.
@@ IV-B Multimodal and Multitask Learning
Recent research attention in multimodal MoE continues to refine interaction-aware routing and modality-specific expert specialization to handle richer and more heterogeneous data: I2MoE xin2025i2moe proposes an interpretable interaction-aware routing framework where distinct experts are trained to capture diverse multimodal interaction patterns, providing local and global interpretability of expert contributions. This strategy enhances modality fusion by explicitly modeling how different sources of information interact during routing. Uni3D-MoE zhang2025uni3dmoe targets 3D multimodal scene understanding by integrating modalities such as multi-view RGB, depth maps, and point clouds. Its learned routing mechanism selects experts based on modality preferences and task-specific context, enabling flexible collaboration for tasks like 3D object recognition and spatial understanding. SMAR xia2025smar introduces a soft modality-aware routing strategy for multimodal MoE models that regularizes routing probabilities across modalities, preserving language capabilities when adapting pretrained MoE models to multimodal objectives. This approach balances modality separation and generalization without altering model architecture or requiring additional modalities during pretraining. MoDES huang2025modes focuses on efficient multimodal MoE inference, proposing a training-free dynamic expert skipping mechanism tailored for heterogeneous contributions across experts and layers in multimodal large language models, decreasing computation while preserving accuracy.
@@ IV-C Healthcare and Life Sciences
Recent work after May 2025 further explores MoE for robust multimodal clinical prediction and anatomically informed medical imaging. MoE-Health wang2025moe_health introduces a dynamic gating Mixture-of-Experts framework tailored for healthcare prediction with heterogeneous and incomplete input modalities (electronic health records, clinical notes, and medical imaging). MoE-Health dynamically selects and integrates specialized expert modules based on available data, improving performance on tasks such as in-hospital mortality, length-of-stay estimation, and readmission prediction under real-world variability in data availability. Such robustness is crucial for clinical deployment where complete multimodal inputs are rarely guaranteed. In medical imaging, MedMoE chopra2025medmoe proposes modality-specialized expert branches within a vision-language MoE framework that adapts feature extraction to diagnostic context, enhancing alignment between imaging and textual findings across diverse clinical benchmarks. Additionally, REN (Regional Expert Networks) peltekian2025ren leverages anatomical priors to train region-specific expert modules for interstitial lung disease diagnosis, demonstrating improved classification performance and interpretability by coupling expert outputs with anatomical structure information. A related trend involves specialized MoE strategies that improve robustness and generalization in medical multimodal learning. For example, adaptive expert grouping mechanisms—designed to exploit collaborative tendencies among experts—have been introduced to reduce routing overhead while preserving generalization benefits in medical contexts ICLR2026FineGrainedMedicalMoE.
A significant frontier remains embodied intelligence in healthcare, where robotic and assistive systems integrate perception, planning, and action in physical settings. As shown in Figure 8, embodied agents leverage cross-modal perception and learned policies across tasks ranging from bedside assistance to rehabilitation. However, barriers such as integration into clinical workflows, simulation-to-reality gaps, and the absence of standardized evaluation benchmarks persist, limiting widespread deployment. Efforts that tightly couple MoE decision mechanisms with domain constraints and real-world validation pipelines may help bridge this gap.
@@ IV-E Natural Language Processing and Large Language Models
Recent work continues advancing MoE design for NLP. MoxE introduces extended LSTM-based MoE with entropy-aware routing to balance expert utilization and improve efficiency in language modeling, showing improved scalability with rare token handling moxe2025. Another direction is lightweight low-rank adaptation Mixture-of-Experts (L-MoE), which unifies MoE and LoRA in an end-to-end trainable framework where task-specialized low-rank expert adapters are dynamically composed via differentiable routing, improving parameter efficiency and dynamic skill composition lmoe2025.
@@ IV-F Methodological Innovation and Theoretical Foundations
Recent work further strengthens the theoretical foundation of MoE systems by reframing routing and expert interaction as structured optimization problems. Omi et al. omi2025similarity_router introduce a similarity-preserving load-balancing objective that enforces consistent expert assignment for related inputs, yielding provable reductions in expert collapse and variance over long training horizons. Complementarily, MaxScore routing dong2025maxscore formulates expert selection as a minimum-cost maximum-flow problem, providing an explicit optimization framework that trades off expert capacity, token assignment, and communication cost. These formulations move beyond heuristic regularization and offer principled guarantees on stability and efficiency.